find the slope of the line through each pair of points
(7,6) (-4,5)
1
Expert's answer
2013-02-07T08:33:20-0500
Let's find the equation of the line passing through this points. The common equation for the line passing through the points (x1,y1) and (x1,y1) is

(x-x2)/(x1-x2) = (y-y2)/(y1-y2).

In our case,

(x-(-4))/(7-(-4)) = (y-5)/(6-5),

or, equally,

(x+4)/11 = y-5.

So, the equation for the line is

y = x/11 + 59/11,

therefore the slope of the line is 1/11.